OpenStudy (anonymous):

Whats the sum of the first 30 terms of the sequence below. a^n=4n+1

OpenStudy (anonymous):

\[\sum_{n = 0}^{29}(4n + 1) = 30 + 4\sum_{n-1}^{29}n\]= 30 + 4(29)(30)/2 = 1770.